OpenStudy (anonymous):

Express each of the following in terms of sinθ and/or cosθ. a) cosθ1/sec

OpenStudy (anonymous):

is this \[\cos(x)\times \frac{1}{\sec(x)}\] ?

OpenStudy (anonymous):

it is unclear from what you wrote, but if that is it, secant is the reciprocal of cosine, so your answer would be \[\cos(x)\times \cos(x)=\cos^2(x)\]